# 1NLPR, MAIS, CASIA 2UCAS 3PKU 4WHU 5ByteDance

### v2

@@ D.2 Qualitative Results on VideoRefer-Bench
Detailed Localized Video Captioning. In Figure 9, we provide qualitative results of extending GAR-8B to generate detailed video descriptions on VideoRefer-Bench [61]. In most cases, where videos usually remain static, GAR-8B manages to generate detailed, specific, and precise descriptions. However, as demonstrated in the last example, GAR-8B fails to capture detailed temporal differences among frames, leading to a low score on “temporal description”. This is because our GAR models are actually trained with only images and lack fine-grained temporal comprehension capabilities.
Detailed Video Understanding. In Figure 10, we provide qualitative results of extending GAR-8B to detailed video understanding on VideoRefer-Bench. GAR-8B is capable of understanding basic motions under a zero-shot setting, e.g., the sequential question, the relation question, and the reasoning question. However, on the “future prediction” protocol, GAR-8B sometimes fails to choose correctly with significant motion changes.
@@ Appendix E Limitation and Failure Cases
One potential limitation is that our GAR is limited to static images. Although it can be successfully extended to video and even achieves competitive results compared with video models (please refer to Tables 6 and 7 for detailed experimental results), it sometimes fails when input videos contain significant motion changes. Specifically, as demonstrated in the failure cases in Figures 9 and 10, GAR-8B is superior at comprehending and describing static videos, and is also capable of understanding basic motions. However, with significant motion changes, GAR-8B sometimes fails. Carefully collecting video training data is a potential solution.
@@ Appendix F Discussion on DLC-Bench
Our analysis in Figure 11 reveals a significant weakness in the original judger of DLC-Bench [22], which relies on a text-only LLM, i.e., LLaMA3.1-8B [13], for automated scoring. Specifically, a fundamental flaw in the original DLC-Bench [22] evaluation lies in its assumption that semantic categories can be accurately adjudicated within the abstract confines of language space alone. However, Figure 11 demonstrates that this text-only approach is inherently unreliable due to the ambiguity of linguistic labels without visual contexts. We argue that the image is the only ground truth capable of resolving this ambiguity. Therefore, we provide the image for valid evaluation. To truly assess a model’s descriptive power, the judge must be multimodal, capable of grounding the generated caption in the visual reality it purports to describe.

### v3

@@ D.2 Qualitative Results on VideoRefer-Bench
Detailed Localized Video Captioning. In Figure 10, we provide qualitative results of extending GAR-8B to generate detailed video descriptions on VideoRefer-Bench (Yuan et al., 2025b). In most cases, where videos usually remain static, GAR-8B manages to generate detailed, specific, and precise descriptions. However, as demonstrated in the last example, GAR-8B fails to capture detailed temporal differences among frames, leading to a low score on “temporal description”. This is because our GAR models are actually trained with only images and lack fine-grained temporal comprehension capabilities.
Detailed Video Understanding. In Figure 11, we provide qualitative results of extending GAR-8B to detailed video understanding on VideoRefer-Bench. GAR-8B is capable of understanding basic motions under a zero-shot setting, e.g., the sequential question, the relation question, and the reasoning question. However, on the “future prediction” protocol, GAR-8B sometimes fails to choose the best choice with significant motion changes.
@@ Appendix E Limitation and Failure Cases
One potential limitation is that our GAR is limited to static images. Although it can be successfully extended to video and even achieves competitive results compared with video models (please refer to Tables 6 and 7 for detailed experimental results), it sometimes fails when input videos contain significant motion changes. Specifically, as demonstrated in the failure cases in Figures 10 and 11, GAR-8B is superior at comprehending and describing static videos, and is also capable of understanding basic motions. However, with significant motion changes, GAR-8B sometimes fails. Carefully collecting video training data could be a potential solution.
@@ Appendix F Discussion on DLC-Bench
Our analysis in Figure 12 reveals a significant weakness in the original judger of DLC-Bench (Lian et al., 2025), which relies on a text-only LLM, i.e., LLaMA3.1-8B (Grattafiori et al., 2024), for automated scoring. Specifically, a fundamental flaw in the original DLC-Bench (Lian et al., 2025) evaluation lies in its assumption that semantic categories can be accurately adjudicated within the abstract confines of language space alone. However, Figure 12 demonstrates that this text-only approach is inherently unreliable due to the ambiguity of linguistic labels without visual contexts. We argue that the image is the only ground truth capable of resolving this ambiguity. Therefore, we provide the image for valid evaluation. To truly assess a model’s descriptive power, the judge must be multimodal, capable of grounding the generated caption in the visual reality it purports to describe.